What is the pH of a 0.02M solution of potassium benzoate solution ( K+ -OOCC6H5). The pka of benzoic acid is 4.19. (Write answer to the hundredths place)
Answer #1

use:

pKa = -log Ka

4.19 = -log Ka

Ka = 6.457*10^-5

use:

Kb = Kw/Ka

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

Kb = (1.0*10^-14)/Ka

Kb = (1.0*10^-14)/6.457*10^-5

Kb = 1.549*10^-10

C6H5COO- dissociates as

C6H5COO- + H2O -----> C6H5COOH + OH-

0.02 0 0

0.02-x x x

Kb = [C6H5COOH][OH-]/[C6H5COO-]

Kb = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Kb = x*x/(c)

so, x = sqrt (Kb*c)

x = sqrt ((1.549*10^-10)*2*10^-2) = 1.76*10^-6

since c is much greater than x, our assumption is correct

so, x = 1.76*10^-6 M

use:

pOH = -log [OH-]

= -log (1.76*10^-6)

= 5.7545

use:

PH = 14 - pOH

= 14 - 5.7545

= 8.2455

Answer: 8.25